How many moles are in 615.90g of Al2(SO4)3


1
Expert's answer
2021-05-04T14:22:02-0400

The number of moles can be calculated as:

n = m / Mr

where n - number of moles, m - mass, Mr - molar mass.

As the molar mass of Al2(SO4)3 equals 342.15 g/mol, the number of moles is:

n = m / Mr = 615.90 g / 342 g/mol = 1.80 mol


Answer: 1.80 mol